# GP+: A Python Library for Kernel-based learning via Gaussian Processes

### v1

@@ 4 Kernel-based Learning
The vanilla GP formulations reviewed in Section 2 break down in high dimensions or with large samples, do not directly accommodate MF modeling or MFBO, and cannot handle categorical features. Since the scalability issue of GPs is rigorously studied in many recent works, in GP+ we focus on holistically addressing the latter two limitations based on the ideas that were first introduced in [71]. In particular, we generalize the concept of kernel-based learning for GPs by introducing new bases and kernels with customized parametric functions that directly enable probabilistic learning from multi-source data and handling qualitative features. Compared to existing works that also develop new kernels for GPs (see for example [72, 73, 74] for handling categorical inputs, building multi-response emulators, and MF modeling), our functions are quite versatile and produce nonlinearly learned embeddings that, while being low-dimensional and highly interpretable, enable GPs to model more complex relations.
To explain our kernel-based approach, we consider an emulation scenario where the input space includes two qualitative features $t_{1}=\{Math,Chemistry\}$ and $t_{2}=\{Japan,France,Canada\}$ which have $l_{1}=2$ and $l_{2}=3$ levels, respectively. Vanilla GPs cannot directly work with $\boldsymbol{t}=[t_{1},t_{2}]^{T}$ since typical kernels such as those in Section 2 require each feature to be associated with a distance metric while categorical variables naturally lack such measures. As schematically illustrated in Figure 2, we address this limitation by first endowing the categorical variables $\boldsymbol{t}=[t_{1},\cdots,t_{dt}]^{T}$ with the quantitative prior representations $\boldsymbol{\pi}_{t}=f_{\pi}(\boldsymbol{t})$ where $f_{\pi}(\cdot)$ is a deterministic user-specified function. These priors are typically high dimensional (i.e., $d\pi>dt$) and can be designed in many ways (we describe some of these below, see the documentation of GP+  for more options). To reduce the dimensionality of these representations while learning the effects of $\boldsymbol{t}$ on the response, we then pass $\boldsymbol{\pi}_{t}$ through the parametric embedding function $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ to obtain $\boldsymbol{h}$ which is a $dh$ dimensional latent representation of $\boldsymbol{t}$ where $d\pi\gg dh$ Since $\boldsymbol{h}=f_{h}(f_{\pi}(\boldsymbol{t});\boldsymbol{\theta}_{h})$ are quantitative, they can be easily used to develop new kernels. For instance, we can extend the Gaussian and Matérn correlation functions as:
@@
In Figure 2 we schematically demonstrate a few options for designing $\boldsymbol{\pi}_{t}$ and the embedding functions for an emulation example where the feature space has quantitative variables $\boldsymbol{x}$ and the two categorical variables $\boldsymbol{t}=[t_{1},t_{2}]$ mentioned above. As shown in the top row of the embedding block, $f_{\pi}(\cdot)$ can simply be a deterministic bijective [^9] Surjective and injective functions may also be used especially if some prior knowledge encourages such choices. We focus on bijective functions in this paper and leave other choices for future studies. function that $(1)$ groups the one-hot-encoded representations of $\boldsymbol{t}$ into a single matrix (this option is the default in GP+), $(2)$ builds a random matrix whose unique rows correspond to the unique combinations of $\boldsymbol{t}$, or $(3)$ constructs multiple matrices where each one corresponds to the one-hot-encoding of one of the categorical variables. The second row in the embedding block of Figure 2 illustrates two options for $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ that consist of parametric matrices (denoted by $\boldsymbol{A}$ in Figure 2) and feed forward neural networks (FFNNs). The construction of $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ is affected by $f_{\pi}(\cdot)$. For instance, the row size of $\boldsymbol{A}$ depends on $\boldsymbol{\pi}_{t}$ [^10] Hence, the random prior encoding can work with a smaller $\boldsymbol{A}$ compared to the grouped one-hot-encoding. while its column size is chosen by the user and determines the dimensionality of the to-be-learnt embedding. The number of $\boldsymbol{A}$ matrices also depends on the priors since it should match with the number of matrices that $f_{\pi}(\cdot)$ generates. We note that these dependencies are automatically enforced in GP+ and the available options for $f_{\pi}(\cdot)$ or $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ in GP+ can be easily accessed by changing its default settings.
Once the parametrized embedding are constructed with any of the procedures described above (or other settings available in GP+), they are concatenated with the numerical features and used in our reformulated mean and covariance functions to build the likelihood function. Then, all the model parameters are estimated via MAP. In the following subsections, we elaborate on how these embeddings as well as reformulated mean and covariance functions benefit MF modeling, inverse parameter estimation, and MFBO.
@@ 4.1 Multi-fidelity Modeling via Deterministic Embedding
The premise of MF modeling is to leverage low-fidelity (LF) data to reduce the reliance on expensive high-fidelity (HF) samples in many-query applications such as design optimization [75, 76], uncertainty propagation and variance reduction [77], BO, calibration of computer models [78], and sensitivity analysis [79]. With the exception of a few recent works such as [80, 81, 3], most existing MF techniques fuse only two data sources while imposing a specific functional relation between them. For instance, the method of Kennedy and O’Hagan (KOH) [22] and its various extensions [82, 19, 18, 21, 83, 84, 85, 20, 86] fuse the HF and LF data based on the following generic relation:
@@
where $y_{h}(\boldsymbol{x})$ and $y_{l}(\boldsymbol{x},\boldsymbol{\zeta})$ denote the HF and LF sources, respectively, $y_{b}(\boldsymbol{x},\boldsymbol{\zeta})$ is the bias function that aims to quantify the systematic bias of the LF data source, $\boldsymbol{\zeta}$ are the calibration parameters whose values must be inversely estimated during the fusion process (see Section 4.4), and $\varepsilon$ denotes normal noise whose variance may be known or not. Equation 14 is based on some rather strong assumptions that do not always hold in practice (e.g., existence of only one LF data source whose bias is additive). To dispense with such inflexible assumptions, the MF modeling capabilities of GP+ are based on converting the fusion process into a nonlinear latent variable learning problem [26].
@@
Figure 3: Graphical representation of multi-fidelity modeling techniques: The method of KOH [22] (a) and its extension to hierarchical techniques (b) impose specific relations between the data sources. However, GP+ (c) does not impose any prior relation among the data sources and its structure resembles an undirected graph.
@@
We highlight that these learnt latent distances provide an average measure of correlation among the sources and cannot identify local discrepancies since the encodings in the $z-$space are not functions of $\boldsymbol{x}$ or $\boldsymbol{t}$. We also note that MF modeling in GP+ differs significantly from most existing methods in that its structure does not prioritize learning any source (e.g., the HF source) over the others. Rather, it aims to integrate all the data sets together to improve its accuracy in emulating all the sources. In this paper, we do not explore the possibility of prioritizing emulation of a particular source (e.g., the HF source) but note that this direction can be pursued in a number of ways such as constraining the embeddings, penalizing the objective function in Equation 8, or designing specific priors.
@@ 4.1.1 Source-Dependent Noise Modeling
Noise inevitably arises in most applications and incorrectly modeling it reduces the performance of any emulator. As mentioned in Section 2, GPs model noise via the nugget or jitter parameter, $\delta$, which changes the covariance matrix from $\boldsymbol{C}$ to $\boldsymbol{C}_{\delta}=\boldsymbol{C}+\delta\boldsymbol{I}_{nn}$. Although this approach works quite well in SF problems, it does not yield the same benefits in MF emulation due to the dissimilar nature of the data sources and their corresponding noises. Consider a bi-fidelity scenario where the HF data comes from an experimental setup and is subject to measurement noise, while the LF data is generated by a deterministic computer code that has a systematic bias due to missing physics. In this case, using only one nugget parameter for MF emulation is obviously not an optimum choice.
To address this issue effectively, we follow [87] and use a nugget vector $\boldsymbol{\delta}=[\delta_{1},\delta_{2},\dots,\delta_{ds}]$ to modify the covariance matrix:
@@ 4.2 Multi-fidelity Modeling via Probabilistic Embedding
To obtain the probabilistic latent representation of the categorical source indicator variable $s$, we reformulate $f_{z}(\boldsymbol{\pi}_{s};\boldsymbol{\theta}_{z})$ to obtain the conditional distribution $q(\boldsymbol{z}\mid s)$. To this end we use the reparameterization trick [88] and design $f_{z}(\boldsymbol{\pi}_{s};\boldsymbol{\theta}_{z})$ accordingly. Specifically, we model $q(\boldsymbol{z}\mid s)$ via a multi-variate normal distribution [^11] Other distributions can also be used but we have had great success with simple ones such as the bivariate normal distribution when $dz=2$. that is fully characterized via the mean vector $\boldsymbol{\mu}_{z}$ and covariance matrix $\boldsymbol{\Sigma}_{z}=\boldsymbol{L}_{z}\boldsymbol{L}_{z}^{T}$ where $\boldsymbol{L}_{z}$ denotes the lower Cholesky decomposition of $\boldsymbol{\Sigma}_{z}$. $\boldsymbol{z}$ is then obtained by:
@@
We note that GP+ by default only builds a probabilistic encoding for $s$ and not $\boldsymbol{t}$ to avoid overfitting: since the number of categorical variables and their levels is typically much larger than the number of levels of $s$ (which indicates the number of sources), probabilistically encoding $\boldsymbol{t}$ requires an FFNN with a large number of parameters and hence may result into overfitting especially if the training data is small. An alternative approach (which can be easily achieved in GP+ by changing its default parameters) is to encode a subset of $\boldsymbol{t}$ in a probabilistic latent space. We leave pursuing this direction to our future studies as it is application specific.
Figure 4: Probabilistic multi-fidelity modeling via GP+: Categorical inputs $\boldsymbol{t}$ are mapped to latent points in the $h-$space while the source indicator variable $s$ is mapped to a conditional distribution in $z-$space. Both mappings are achieved via deterministic and differentiable functions. Due to the probabilistic nature of $\boldsymbol{z}$, multiple forward passes are required to obtain the final outputs of the model.
With a probabilistic fidelity embedding, we must sample from $q(\boldsymbol{z}\mid s)$ multiple times during both training and testing since even for fixed $s,\boldsymbol{t},$ and $\boldsymbol{x}$ the predicted covariance from the GP model varies due to $\boldsymbol{\epsilon}_{z}$ (the effect on the mean function depends on its formulation, see Section 4.3). Hence, for any fixed values of $s,\boldsymbol{t},\boldsymbol{x},$ and model parameters (i.e., $\boldsymbol{\beta},\boldsymbol{\theta}_{h},\sigma^{2},$ and $\boldsymbol{\theta}_{z}$), we generate $M$ samples from $q(\boldsymbol{z}\mid s)$ [^12] These samples are generated by drawing $M$ random $\boldsymbol{\epsilon}_{z}$ vectors of size $dz$ from a standard multi-variate normal distribution. to build an ensemble or mixture of $M$ GPs. Since $\boldsymbol{\epsilon}_{z}$ are independent and identically distributed (IID), each member of the GP ensemble is equally probable and we can extend Equation 10b via the laws of total expectation and covariance [89, 90]. To this end, we must obtain expressions for $\boldsymbol{m}$ and $\boldsymbol{C}$ in Equation 8 during training. We distinguish among the GPs in the mixture model via the random variable $I$ (whose probability mass function or PMF is $p(I=k)=\frac{1}{M}$ for $k=1,\cdots,M$) and we calculate the ensemble mean as:
@@
Inserting Equations 29 and 30 into Equation 28 we obtain:
@@
For prediction, we take $Q$ samples from the probabilistic fidelity embedding to determine the ensemble mean and variance:
@@
where $\mu_{k}(\boldsymbol{u}^{*})$ and $\tau_{k}^{2}(\boldsymbol{u}^{*})$ are the mean and variance of the $k^{th}$ GP and obtained via Equation 9b. We note that the ensemble variance formula is a specific instance of the ensemble covariance given in Equation 31. Additionally, we typically use $Q>M$ to reduce the training costs.

### v2

@@ 4 Kernel-based Learning
The vanilla GP formulations reviewed in Section 2 break down in high dimensions or with large samples [hensman2013gaussian, bengio2005curse], do not directly accommodate MF modeling or MFBO [eweis2022data, foumani2023multi], and cannot handle categorical features [oune2021latent]. Since the scalability issue of GPs is rigorously studied in many recent works, in GP+ we focus on holistically addressing the latter two limitations based on the ideas that were first introduced in [oune2021latent]. In particular, we generalize the concept of kernel-based learning for GPs by introducing new bases and kernels with customized parametric functions that directly enable probabilistic learning from multi-source data and handling qualitative features. Compablack to existing works that also develop new kernels for GPs (see for example [tao2021latent, bonilla2007multi, poloczek2017multi] for handling categorical inputs, building multi-response emulators, and MF modeling), our functions are quite versatile and produce nonlinearly learned embeddings that, while being low-dimensional and highly interpretable, enable GPs to model more complex relations.
To explain our kernel-based approach, we consider an emulation scenario where the input space includes two qualitative features $t_{1}=\{Math,Chemistry\}$ and $t_{2}=\{Japan,France,Canada\}$ which have $l_{1}=2$ and $l_{2}=3$ levels, respectively. Vanilla GPs cannot directly work with $\boldsymbol{t}=[t_{1},t_{2}]^{T}$ since typical kernels such as those in Equation 4c require each feature to be associated with a distance metric while categorical variables naturally lack such measures. As schematically illustrated in Figure 2, we address this limitation by first endowing the categorical variables $\boldsymbol{t}=[t_{1},\cdots,t_{dt}]^{T}$ with the quantitative prior representations $\boldsymbol{\pi}_{t}=f_{\pi}(\boldsymbol{t})$ where $f_{\pi}(\cdot)$ is a deterministic user-specified function. These priors are typically high dimensional (i.e., $d\pi>dt$) and can be designed in many ways (we describe some of these below, see supplementary comments on our GitHub page for more options). To blackuce the dimensionality of these representations while learning the effects of $\boldsymbol{t}$ on the response, we then pass $\boldsymbol{\pi}_{t}$ through the parametric embedding function $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ to obtain $\boldsymbol{h}$ which is a $dh$ dimensional latent representation of $\boldsymbol{t}$ where $d\pi\gg dh$. Since $\boldsymbol{h}=f_{h}(f_{\pi}(\boldsymbol{t});\boldsymbol{\theta}_{h})$ are quantitative, they can be easily used to develop new kernels. For instance, we can extend the Gaussian and Matérn correlation functions as:
@@
In Figure 2 we schematically demonstrate a few options for designing $\boldsymbol{\pi}_{t}$ and the embedding functions for an emulation example where the feature space has quantitative variables $\boldsymbol{x}$ and the two categorical variables $\boldsymbol{t}=[t_{1},t_{2}]$ mentioned above. As shown in the top row of the embedding block, $f_{\pi}(\cdot)$ can simply be a deterministic bijective [^10] Surjective and injective functions may also be used especially if some prior knowledge encourages such choices. We focus on bijective functions in this paper and leave other choices for future studies. function that $(1)$ groups the one-hot-encoded representations of $\boldsymbol{t}$ into a single matrix (this option is the default in GP+), $(2)$ builds a random matrix whose unique rows correspond to the unique combinations of $\boldsymbol{t}$, or $(3)$ constructs multiple matrices where each one corresponds to the one-hot-encoding of one of the categorical variables. The second row in the embedding block of Figure 2 illustrates two options for $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ that consist of parametric matrices (denoted by $\boldsymbol{A}$ in Figure 2) and FFNNs. The construction of $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ is affected by $f_{\pi}(\cdot)$. For instance, the row size of $\boldsymbol{A}$ depends on $\boldsymbol{\pi}_{t}$ [^11] Hence, the random prior encoding can work with a smaller $\boldsymbol{A}$ compablack to the grouped one-hot-encoding. while its column size is chosen by the user and determines the dimensionality of the to-be-learnt embedding. The number of $\boldsymbol{A}$ matrices also depends on the priors since it should match with the number of matrices that $f_{\pi}(\cdot)$ generates. We note that these dependencies are automatically enforced in GP+ and the available options for $f_{\pi}(\cdot)$ or $f_{h}(\boldsymbol{\pi}_{t};\boldsymbol{\theta}_{h})$ in GP+ can be easily accessed by changing its default settings.
Once the parametrized embedding is constructed with any of the procedures described above (or other settings available in GP+), they are concatenated with the numerical features and used in our reformulated mean and covariance functions to build the likelihood function. Then, all the model parameters are estimated via MAP. In the following subsections, we elaborate on how these embeddings as well as reformulated mean and covariance functions benefit MF modeling, inverse parameter estimation, and MFBO.
@@ 4.1 Multi-fidelity Modeling via Deterministic Embedding
The premise of MF modeling is to leverage low-fidelity (LF) data to blackuce the reliance on expensive high-fidelity (HF) samples in many-query applications such as design optimization [chakraborty2017surrogate, korondi2021multi], uncertainty propagation and variance blackuction [dixon2023covariance], BO, calibration of computer models [absi2016multi], and sensitivity analysis [RN452]. With the exception of a few recent works such as [RN1770, RN1876, RN1845], most existing MF techniques fuse only two data sources while imposing a specific functional relation between them. For instance, the method of Kennedy and O’Hagan (KOH) [RN705] and its various extensions [RN648, RN272, RN271, RN337, RN344, RN335, RN336, RN334, RN270] fuse the HF and LF data based on the following generic relation:
@@
where $y_{h}(\boldsymbol{x})$ and $y_{l}(\boldsymbol{x},\boldsymbol{\zeta})$ denote the HF and LF sources, respectively, $y_{b}(\boldsymbol{x},\boldsymbol{\zeta})$ is the bias function that aims to quantify the systematic bias of the LF data source, $\boldsymbol{\zeta}$ are the calibration parameters whose values must be inversely estimated during the fusion process (see Section 4.4), and $\varepsilon$ denotes normal noise whose variance may be known or not. Equation 14 is based on some strong assumptions that do not always hold in practice (e.g., existence of only one LF data source whose bias is additive). To dispense with such inflexible assumptions, the MF modeling capabilities of GP+ are based on converting the fusion process into a nonlinear latent variable learning problem [RN1838].
@@
Figure 3: Graphical representation of multi-fidelity modeling techniques: The method of KOH [RN705] (a) and its extension to hierarchical techniques (b) impose specific relations between the data sources. However, GP+ (c) does not impose any prior relation among the data sources and its structure resembles an undirected graph.
@@
We highlight that these learned latent distances provide an average measure of correlation among the sources and cannot identify local discrepancies since the encodings in the $z-$space are not functions of $\boldsymbol{x}$ or $\boldsymbol{t}$. Thus, if some LF sources are only locally correlated with the HF source, they will be encoded relatively far from the HF source in the learned embedding. This implies that those sources provide valuable insights in certain areas of the domain and keeping or dropping them depends on the specific applications. For instance, if the application is emulation, techniques such as cross-validation or train-test splits can assist in determining which sources to keep or drop. However if the goal is multi-fidelity BO (which typically starts with very small initial data, especially from the HF source), we recommend keeping all the sources during the optimization process (see [foumani2023effects] for more details).
MF modeling in GP+ differs significantly from most existing methods in that its structure does not prioritize learning any source (e.g., the HF source) over the others, i.e., GP+ aims to integrate all the data sets together to improve its accuracy in emulating all the sources. For example, multilevel best linear unbiased estimators (MBLUE) and approximate control variate (ACV) are two variance blackuction-based techniques that leverage MF data to more accurately learn the HF source [MLBLUE, ACV]. Since these methods prioritize surrogating the HF source and do not build surrogates for the LF sources, they cannot be used in applications such as MFBO where one has to emulate all sources. In this paper, we do not explore the possibility of prioritizing emulation of a particular source (e.g., the HF source) but note that this direction can be pursued in a number of ways such as constraining the embeddings, penalizing the objective function in Equation 8, or designing specific priors.
@@ 4.1.1 Source-Dependent Noise Modeling
Noise inevitably arises in most applications and incorrectly modeling it blackuces the performance of any emulator. As mentioned in Section 2, GPs model noise via the nugget or jitter parameter, $\delta$, which changes the covariance matrix from $\boldsymbol{C}$ to $\boldsymbol{C}_{\delta}=\boldsymbol{C}+\delta\boldsymbol{I}_{nn}$. Although this approach works quite well in SF problems, it does not yield the same benefits in MF emulation due to the dissimilar nature of the data sources and their corresponding noises. Consider a bi-fidelity scenario where the HF data comes from an experimental setup and is subject to measurement noise, while the LF data is generated by a deterministic computer code that has a systematic bias due to missing physics. In this case, using only one nugget parameter for MF emulation is obviously not an optimum choice.
To address this issue effectively, we follow [foumani2023effects] and use a nugget vector $\boldsymbol{\delta}=[\delta_{1},\delta_{2},\dots,\delta_{ds}]$ to modify the covariance matrix:
@@
We highlight that the above formulations model stationary noise for each data source. We make this choice due to the fact that modeling an input-dependent noise increases the number of hyperparameters by at least $ds\times dx$ which can result in overfitting [RN340, RN717]. Therefore, to balance the risk of overfitting with the uncertainty quantification capacity of our emulator, we assume the noise variance is not a function of the input variables and only depends on the data source.
@@ 4.2 Multi-fidelity Modeling via Probabilistic Embedding
To obtain the probabilistic latent representation of the categorical source indicator variable $s$, we reformulate $f_{z}(\boldsymbol{\pi}_{s};\boldsymbol{\theta}_{z})$ to obtain the conditional distribution $q(\boldsymbol{z}\mid s)$. To this end we use the reparameterization trick [kingma2013auto] and design $f_{z}(\boldsymbol{\pi}_{s};\boldsymbol{\theta}_{z})$ accordingly. Specifically, we model $q(\boldsymbol{z}\mid s)$ via a multi-variate normal distribution [^12] Other distributions can also be used but we have had great success with simple ones such as the bivariate normal distribution when $dz=2$. that is fully characterized via the mean vector $\boldsymbol{\mu}_{z}$ and covariance matrix $\boldsymbol{\Sigma}_{z}=\boldsymbol{L}_{z}\boldsymbol{L}_{z}^{T}$ where $\boldsymbol{L}_{z}$ denotes the lower Cholesky decomposition of $\boldsymbol{\Sigma}_{z}$. $\boldsymbol{z}$ is then obtained by:
@@
We note that GP+ by default only builds a probabilistic encoding for $s$ and not $\boldsymbol{t}$ to avoid overfitting: since the number of categorical variables and their levels is typically much larger than the number of levels of $s$ (which indicates the number of sources), probabilistically encoding $\boldsymbol{t}$ requires an FFNN with a large number of parameters and hence may result into overfitting especially if the training data is small. An alternative approach (which can be achieved in GP+ by changing its default parameters) is to encode a subset of $\boldsymbol{t}$ in a probabilistic latent space. We leave pursuing this direction to our future studies as it is application specific.
Figure 4: Probabilistic multi-fidelity modeling via GP+: Categorical inputs $\boldsymbol{t}$ are mapped to latent points in the $h-$space while the source indicator variable $s$ is mapped to a conditional distribution in $z-$space. Both mappings are achieved via deterministic and differentiable functions. Due to the probabilistic nature of $\boldsymbol{z}$, multiple forward passes are requiblack to obtain the final outputs of the model.
[…]
@@
Inserting Equations 23 and 24 into Equation 22 we obtain:
@@
For pblackiction, we take $Q$ samples from the probabilistic fidelity embedding to determine the ensemble mean and variance:
@@
where $\mu_{k}(\boldsymbol{u}^{*})$ and $\tau_{k}^{2}(\boldsymbol{u}^{*})$ are the mean and variance of the $k^{th}$ GP and obtained via Equation 9b. We note that the ensemble variance formula is a specific instance of the ensemble covariance given in Equation 25. Additionally, we typically use $Q>M$ to blackuce the training costs.